A news organization surveyed 80 adults. Each said he or she gets news from only one source. Here is a summary of their sources of news.
Source of news Number of adults
Internet 21
Newspaper 23
Radio 13
Television 23
Three of the adults from the survey are selected at random, one at a time without replacement. What is the probability that none of the three adults get news from the newspaper?
Do not round your intermediate computations. Round your final answer to three decimal places.
1
Expert's answer
2020-09-29T18:56:54-0400

The probability that an adult does not get news from a newspaper is: "\frac{23}{80}" .

The probability that the second selected adult also does not get news from a newspaper is: "\frac{23}{79}" . For the third one we get: "\frac{23}{78}" .

Thus, we receive: "\frac{23}{80}\frac{23}{79}\frac{23}{78}=\frac{12167}{492960}\approx0.025"

Answer:0.025 (rounded to 3 decimal places)